OpenStudy (anonymous):

**PLEASE HELP! FAN AND MEDAL!!!** 4/5x + 4/3 = 2x solve for X

OpenStudy (anonymous):

4/(5x) + 4/3 = 2x

jigglypuff314 (jigglypuff314):

multiply both sides by (5x)(3) so you get 4(3) + 4(5x) = (2x)(5x)(3) simplify 12 + 20x = 30x^2 bring all the values to one side 30x^2 - 20x - 12 = 0 then factor out greatest common factor and factor.

OpenStudy (campbell_st):

multiply every term by 15x this wll eliminate the denominators in the 2 fractions so you get 12 + 20x = 30x^2 or 30x^2 - 20x -12 = 0 divide by 2 15x^2 - 10x - 6 = 0 now you have a quadratic to solve.. hope this helps

OpenStudy (anonymous):

so I would solve 15x^2 - 10x - 6 = 0?

OpenStudy (campbell_st):

yes.... probably general quadratic formula is best

OpenStudy (anonymous):

what if I need it as a fraction?

OpenStudy (campbell_st):

well I think you'll find its an irrational solution....

OpenStudy (anonymous):

cause my thing says that im supposed to end up with a fraction :/

OpenStudy (campbell_st):

well if you use the general quadratic formula you will get a fraction.... \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] but its irrational

OpenStudy (anonymous):

is there a way to do it with out the general quadric formula? because I don't think I have even heard of that in class yet

OpenStudy (campbell_st):

well look at the discriminant... \[b^2 - 4ac \] its 460 which isn't a perfect square. so the the GQF seems best... an alternative is complete the square...

jigglypuff314 (jigglypuff314):

hmm... I can't seem to factor it either without quadratic formula... try this? http://www.wolframalpha.com/input/?i=15x%5E2+-+10x+-+6+%3D+0

OpenStudy (anonymous):

its a review of 8th grade math...I don't think it can possably be so complicated

OpenStudy (campbell_st):

ok... well perhaps you should ask you teacher, show the teacher what you have done... and if its 8th grade the problem makes more sense as \[\frac{4}{5} x + \frac{4}{3} = 2x\] by subtracting 4/5 x you get \[\frac{4}{3} = \frac{6}{5} x \] just solve for x